OpenStudy (keni):

What is the surface area of the prism? A rectangular prism with length of 7 meters, width of 2 meters and height of 4 meters.

OpenStudy (keni):

What is the height of a rectangular based pyramid with a volume of four hundred fifty cubic meters and base dimensions of fifteen meters and ten meters

OpenStudy (anonymous):

The formula is: 2(lw)+2(wh)+2(lh)

OpenStudy (anonymous):

l=length w=width h=height

OpenStudy (keni):

ok

OpenStudy (keni):

im still confused

OpenStudy (anonymous):

Why?

OpenStudy (keni):

so is 27

OpenStudy (anonymous):

what do you mean? "so is 27"?

OpenStudy (anonymous):

So we know the formula already. 2(7x2)+2(2x4)+2(7x4) 2(14)+2(8)+2(28) 28+16+56=100

OpenStudy (anonymous):

Formulas for a rectangular prism: Volume of Rectangular Prism: V = lwh Surface Area of Rectangular Prism: S = 2(lw + lh + wh) Space Diagonal of Rectangular Prism: (similar to the distance between 2 points) d = √(l2 + w2 + h2) A cube is a special case where l = w = h. So you can find the volume of a cube or surface area of a cube by setting these values equal to each other. Calculations for a rectangular prism: 1. Given the length, width and height find the volume, surface area and diagonal of a rectangular prism h, l and w are known; find V, S and d V = lwh S = 2(lw + lh + wh) d = √(l2 + w2 + h2) 2. Given the surface area, length and width find the height, volume and diagonal of a rectangular prism S, l and w are known; find h, V and d h = (S - 2lw) / (2l + 2w) V = lwh d = √(l2 + w2 + h2) 3. Given the volume, length and width find the height, surface area, and diagonal of a rectangular prism V, l and w are known; find h, S and d h = V / lw S = 2(lw + lh + wh) d = √(l2 + w2 + h2) 4. Given the diagonal, length and width find the height, volume and surface area of a rectangular prism d, l and w are known; find h, V and S h = √(d2 - l2 - w2) V = lwh S = 2(lw + lh + wh)

OpenStudy (anonymous):

(28) + (16) + (56) add it together.

OpenStudy (anonymous):

The rest of you are just stressing this kid out. Lol

OpenStudy (keni):

so is 1 meter

OpenStudy (anonymous):

....

OpenStudy (anonymous):

When a meter multiples another meter how much do you get?

OpenStudy (anonymous):

Example: 2m * 2m = 4m^2

OpenStudy (anonymous):

You got 100. Now input the meters in.

OpenStudy (anonymous):

100m^2 ~

OpenStudy (keni):

idk

OpenStudy (anonymous):

Kid. 100m^2 is the answer....

OpenStudy (keni):

but on this exam only give 4 options 1 meter 3 meters 9 meters 27 meters

OpenStudy (anonymous):

~copy and past the whole question WITH choices.~

OpenStudy (keni):

What is the height of a rectangular based pyramid with a volume of four hundred fifty cubic meters and base dimensions of fifteen meters and ten meters? 1 meter 3 meters 9 meters twenty-seven meters

OpenStudy (anonymous):

...You have two questions... LMAO

OpenStudy (keni):

no wonder

OpenStudy (anonymous):

That's why I put in all possible formulas. Figured he could take the numbers and plug them in

OpenStudy (anonymous):

I don't know why, but I get 3 meters for that question~

OpenStudy (jdoe0001):

so, I see 2 exercises there, one on a prism and another on a pyramid which is this one?

OpenStudy (anonymous):

I think the person is now on the second one.

OpenStudy (keni):

thank you mikeyy so 3 meters

OpenStudy (anonymous):

~Keni, let jdoe help you. I'm not so sure with my answer. I gave it a educated guess.

OpenStudy (keni):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 8 inches and the radius is 7 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi. one hundred seventy-five point eight four square inches three hundred fifty-one point six eight square inches one thousand two hundred thirty point eight eight square inches one thousand four hundred six point seven two square inches

OpenStudy (keni):

can u help me with that one

OpenStudy (anonymous):

Keni, please make new questions! This is the most annoying thing for people who are trying to help you.

OpenStudy (jdoe0001):

if you post anew, you'd get more visibility and we can all help and revise each other :)

OpenStudy (keni):

ok thank you